# Implicit Bias of SignGD and Adam on Multiclass Separable Data

### v1

@@ Appendix D Proof of Adam
Appendix E Normalized $p$-Nrom Steepest Descent
[…]
Consider normalized steepest-descent with respect to the $p$-norm:
${\bm{W}}_{t+1}={\bm{W}}_{t}-\eta_{t}\bm{\Delta}_{t},\qquad\text{where}~{}~{}\bm{\Delta}_{t}:=\arg\max_{\|\bm{\Delta}\|\leq 1}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\bm{\Delta}\rangle\,.$ (28)
Lemma E.1 generalizes Lemma B.1.
Lemma E.1 (${\mathcal{G}}({\bm{W}})$ as proxy to the loss-gradient norm).
Define the margin $\gamma$ with respect to $p$-norm, for $p\geq 1$ and denote $\|\cdot\|_{\star}$ its dual norm. Then, for any ${\bm{W}}$ it holds that
$2B\cdot{\mathcal{G}}({\bm{W}})\geq\|{\nabla\mathcal{L}({\bm{W}})}\|_{\star}\geq\gamma\cdot{\mathcal{G}}({\bm{W}})\,.$
@@ Appendix E Normalized $p$-Nrom Steepest Descent
$\|{\nabla\mathcal{L}({\bm{W}})}\|_{\star}$ $=\max_{\|{\bm{A}}\|\leq 1}\langle\bm{A},-\nabla\mathcal{L}({\bm{W}})\rangle$
$\geq\max_{\|\bm{A}\|\leq 1}\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}$
$\geq\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\cdot\max_{\|\bm{A}\|\leq 1}\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}.$
The upper bound follows because for $p$-norms with $p\geq 1$
$\|\nabla\mathcal{L}({\bm{W}})\|_{\star}\leq{\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}$
and we can use the bound for ${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}$. ∎
A direct consequence of Lemma E.1 is the following.
Lemma E.2.
@@
We replace the term ${\left\|\nabla\mathcal{L}({\bm{W}}_{0})\right\|_{\rm{sum}}}{\left\|{\bm{W}}_{0}-{\bm{W}}\right\|_{\max}}$ in Lemma B.2 with $\lVert\nabla\mathcal{L}({\bm{W}}_{0})\rVert_{*}\lVert{\bm{W}}_{0}-{\bm{W}}\rVert$. The rest proof follows the same steps as Lemma B.2. ∎
Lemma E.3 generalizes Lemma A.3.
Lemma E.3.
For any $\bm{s}\in\Delta^{k-1}$ in the $k$-dimensional simplex, any index $c\in[k]$, and any ${\bm{v}}\in\mathbb{R}^{k}$ it holds:
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}\leq 4\,\|{\bm{v}}\|^{2}\,(1-s_{c})$

### v4

@@ Appendix B Facts about CE loss and Softmax
Appendix C Lemmas on Loss and Proxy Function
Lemma 16 shows that ${\mathcal{G}}({\bm{W}})$ upper and lower bound the dual norm of the loss gradient.
Lemma 16 (${\mathcal{G}}({\bm{W}})$ as proxy to the loss-gradient norm).
Under Assumption 4. For any ${\bm{W}}\in\mathbb{R}^{k\times d}$, it holds that
$2B\cdot{\mathcal{G}}({\bm{W}})\geq\lVert\nabla\mathcal{L}({\bm{W}})\rVert_{*}\geq\gamma\cdot{\mathcal{G}}({\bm{W}})\,.$
@@ Appendix C Lemmas on Loss and Proxy Function
$\lVert\nabla\mathcal{L}({\bm{W}})\rVert_{*}$ $=\max_{\lVert\bm{A}\rVert\leq 1}\langle\bm{A},-\nabla\mathcal{L}({\bm{W}})\rangle$
$\geq\max_{\lVert\bm{A}\rVert\leq 1}\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}$
$\geq\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\cdot\max_{\lVert\bm{A}\rVert\leq 1}\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}.$
Second, for the upper bound, it holds by triangle inequality and relationships (7) that
$\lVert\nabla\mathcal{L}({\bm{W}})\rVert_{*}\leq{\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}\leq\frac{1}{n}\sum_{i\in[n]}{\left\|\nabla\ell_{i}({\bm{W}})\right\|_{\rm{sum}}}\,$,
$\|\bm{e}_{y_{i}}-\mathbb{S}_{y_{i}}({\bm{W}}\bm{h}_{i})\|_{1}=2(1-s_{y_{i}})$
Built upon Lemma 16, we obtain a simple bound on the loss difference at two points.
Lemma 17.
@@
By convexity of $\mathcal{L}$, we have
$\mathcal{L}({\bm{W}}_{0})-\mathcal{L}({\bm{W}})\leq\langle\nabla\mathcal{L}({\bm{W}}_{0}),{\bm{W}}_{0}-{\bm{W}}\rangle\leq\lVert\nabla\mathcal{L}({\bm{W}}_{0})\rVert_{*}\lVert{\bm{W}}_{0}-{\bm{W}}\rVert\leq 2B\lVert{\bm{W}}_{0}-{\bm{W}}\rVert\,$,
where the last inequality is by Lemma 16. Similarly, we can also show that $\mathcal{L}({\bm{W}})-\mathcal{L}({\bm{W}}_{0})\leq 2B\lVert{\bm{W}}_{0}-{\bm{W}}\rVert$. ∎
Lemma 18 shows the close relationships between ${\mathcal{G}}({\bm{W}})$ and $\mathcal{L}({\bm{W}})$. The proxy ${\mathcal{G}}({\bm{W}})$ not only lower bounds $\mathcal{L}({\bm{W}})$, but also upper bounds $\mathcal{L}({\bm{W}})$ up to a factor depending on $\mathcal{L}({\bm{W}})$. Moreover, the rate of convergence $\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}$ depends on the rate of decrease in the loss.
Lemma 18 (${\mathcal{G}}({\bm{W}})$ as proxy to the loss).
Let ${\bm{W}}\in\mathbb{R}^{k\times d}$, we have
  1. (i)
  2. (ii)